# From Memories to Maps: Mechanisms of In-Context Reinforcement Learning in Transformers

### v2

@@ 1 Introduction
Animals can learn efficiently from sparse data, rapidly adapting to new environments with minimal experience. For example, humans can infer underlying structure or learn new concepts from just a handful of examples and mice in maze tasks can identify optimal paths after only a few successful trials [Meister, 2022]. Standard reinforcement learning (RL) algorithms, which typically rely on incremental value updates to shape decision-making, does not capture this rapid learning behavior well [Eckstein et al., 2024]. One explanation is that biological agents possess useful priors shaped by evolution and experience, allowing them to generalize quickly in naturalistic settings. They also rely on episodic memory, the ability to recall specific past experiences to guide decisions in novel situations.
Here, we ask how episodic memory operates not just as storage, but as a computational substrate for rapid learning and decision-making. We train a transformer model to perform in-context reinforcement learning [Lee et al., 2023] on navigation tasks inspired by rodent behavior. In each new environment, the model receives exploratory trajectories as context and infers a goal-directed policy. Transformers are especially relevant not only because of their established capabilities for rapid and flexible in-context learning [Dong et al., 2022, Brown et al., 2020, Lampinen et al., 2024], but also because their key–value memory architecture has been linked to models of episodic memory in the brain [Krotov and Hopfield, 2020, Tyulmankov et al., 2021, Fang et al., 2025]. Understanding the reinforcement learning strategy that emerges in these models can provide new hypotheses for how memory-based computations might support flexible decision-making in new environments.
We focus on two task suites: spatially regular gridworlds and hierarchically structured tree mazes (Fig 1BC). While both require memory to support goal-directed behavior, they differ sharply in geometry: gridworlds are Euclidean and spatially continuous, whereas tree mazes are non-Euclidean and branch-structured. This contrast allows us to evaluate how learned in-context strategies generalize across structural regimes known to challenge standard sequence models– for instance, language models are known to struggle with symbolic reasoning and hierarchical generalization in tree-structured domains [Bogin and Berant, 2022, Ruiz and Nachum, 2021, Keysers et al., 2020].
@@
  • The model learns computations similar to those that support natural cognition: the representation learning strategies are consistent with suggested roles for hippocampus and entorhinal cortex, and memory retrieval patterns at decision time is consistent with hippocampal replay.
  • Mechanistic analysis of the emergent RL strategy shows that the model does not use standard model-free or model-based RL methods.
@@
  • We give descriptions for the in-context RL algorithm that emerged in transformers to solve the two planning tasks. We find strategies tend to rely on intermediate computations cached in memory tokens, demonstrating how episodic memory can be used as a computational workspace.
@@
Figure 1: A transformer is trained to in-context reinforcement learn in diverse planning tasks. A. Diagram of the meta-learning setup. Each task in the training distribution is defined by a state transition function $\mathcal{T}$, a state encoding function $\mathcal{S}$, and a reward function $\mathcal{R}$. For each task, the model is trained via supervision to predict the optimal action from a query state $s_{\text{query}}$, given a context dataset $\mathcal{D}=\{(s_{1},a_{1},s^{\prime}_{1},r_{1}),(s_{2},a_{2},s^{\prime}_{2},r_{2}),\dots\}$ of RL transition tuples sampled from the task. B. Illustration of three training tasks (orange) and one test task (gray) from the gridworld distribution. In each task, the underlying graph structure is fixed (shown in gray), with a reward location (red star) that varies across tasks. Each state is encoded as a random Gaussian vector (bottom). Importantly, state encodings in test tasks are novel. C. As in (B), but for the tree maze distribution. The graph structures in the training set is drawn from probabilistically branching trees, while the test set graph structure is a full binary tree.
2 Related Works
Meta-learning to discover in-context reinforcement learning algorithms In order to develop a model that can in-context reinforcement learn, we use a recently introduced meta-learning framework [Lee et al., 2023]. Meta-learning is concerned with “learning-to-learn”, using prior knowledge learned from previous tasks to support rapid adaptation to new ones [Beck et al., 2023]. That is, the goal is to learn an algorithm $f$ that can be deployed in new tasks. The learning of $f$ is called the outer-loop while $f$ itself is referred to as the inner-loop [Beck et al., 2023]. In many settings where $f$ is a RL algorithm (meta-RL), the outer loop shapes the weights of the network, but learning within each task occurs via activation dynamics– through memory and internal state– rather than parameter updates [Beck et al., 2023, Lampinen et al., 2024]. The resultant $f$ is considered an in-context learning algorithm [Beck et al., 2023, Sandbrink and Summerfield, 2024, Lampinen et al., 2024].
@@ 2 Related Works
Early examples of this approach include RL2, which meta-trains a recurrent neural network using reinforcement learning in the outer loop, such that an in-context RL strategy emerges in the inner loop [Wang et al., 2016, Duan et al., 2016]. Subsequent work extended this approach to include explicit episodic memory mechanisms, combining RNNs with key–value memory architectures [Ritter et al., 2018, 2020, Team et al., 2023]. More recently, Lee et al. [2023] proposed decision-pretrained transformers (DPTs), which use supervised training in the outer loop to induce in-context RL behavior in the inner loop. We adopt DPTs in this work both for practical reasons–their scalability and ease of training–and for scientific ones: the transformer architecture allows us to probe how memory-based computation supports in-context learning.
Meta-learning to describe cognition and neural activity Meta-learning has been proposed as a framework to model both cognitive flexibility and structured learning in neuroscience and psychology [Binz et al., 2024]. In human cognitive tasks, meta-learned models have been used to replicate observed heuristics in decision-making and to account for few-shot generalization [Dasgupta et al., 2020, Binz et al., 2022, Lake and Baroni, 2023]. Meta-RL, in particular, has been used to generate hypotheses about how neural systems implement learning across tasks. Previous studies have used the RL2 framework to show how the outer and inner learning loops can model different areas of the brain, with the prefrontal cortex often playing a key role [Wang et al., 2018, Hattori et al., 2023, Zheng et al., 2025]. Despite these advances, many computational models in neuroscience rely on single-task RL, training agents independently on each task without leveraging prior experience. In contrast, we use meta-RL as a tool for developing flexible in-context learning, without attempting to localize the outer loop to any specific brain region. Our focus is on the computational content of the learned representations and decision-making strategies.
Transformers and episodic memory systems Transformer models process sequences by computing self-attention over key–value pairs, enabling flexible access to information across long contexts [Vaswani et al., 2017]. This key–value structure has led to interpretations of transformers as memory systems [Geva et al., 2020], aligning them with a broader class of models that incorporate explicit memory mechanisms [Graves et al., 2014, Sukhbaatar et al., 2015, Graves et al., 2016, Banino et al., 2020]. These systems separate memory addressing (keys) from memory content (values), enabling high-fidelity storage and targeted retrieval. This architectural separation bears many similarities to theoretical accounts of episodic memory in the brain [Teyler and DiScenna, 1986, Teyler and Rudy, 2007, Gershman et al., 2025]. For example, recent work has formalized connections between key–value architectures and Hopfield networks, a classic model of associative memory in the brain [Krotov and Hopfield, 2020, Ramsauer et al., 2020]. Other studies have proposed biologically grounded implementations of key–value attention mechanisms, further linking transformer-like architectures to neural computation [Bricken and Pehlevan, 2021, Tyulmankov et al., 2021, Whittington et al., 2021, Kozachkov et al., 2023, Fang et al., 2025, Chandra et al., 2025]. Several of these models take direct inspiration from the hippocampus, a brain region widely implicated in episodic memory [Whittington et al., 2021, Fang et al., 2025, Chandra et al., 2025]. Recent experimental work has also identified key–value–like coding patterns in hippocampal activity during episodic memory tasks [Chettih et al., 2024]. Together, these findings suggest that key–value architectures offer a useful computational analogy for episodic memory. In this work, we analyze a meta-trained transformer to examine what kinds of memory-supported strategies emerge when such an architecture is optimized for rapid adaptation.
@@ 3 Experimental Methods
We adopt the decision-pretraining framework introduced by Lee et al. [2023] (Fig. 1A), where the model is trained via supervision to learn an optimal policy given in-context experience. Each task $\tau$ is a Markov Decision Process (MDP) defined by the tuple $\langle\mathcal{S},\mathcal{A},\mathcal{T},\mathcal{R},\rho\rangle$, with state encoding function $\mathcal{S}$, action space $\mathcal{A}$, transition function $\mathcal{T}$, reward function $\mathcal{R}$, and initial state distribution $\rho$. A subset of tasks from this distribution is sampled for meta-training. For each task, the model is provided an in-context dataset $\mathcal{D}$, consisting of a trajectory collected from an exploratory policy interacting with the environment. Each dataset $\mathcal{D}$ consists of interaction tuples $(s,a,s^{\prime},r)$, where each tuple represents a transition experienced by the agent during exploration. The model receives both the in-context dataset $\mathcal{D}$ and a query state $s_{\text{query}}\in\mathcal{S}$. It is then trained to predict the optimal action $a^{*}\sim\pi^{*}(\cdot\mid s_{\text{query}})$ under the task-specific optimal policy $\pi^{*}$.
At test time, the trained model is evaluated on held-out tasks, using only in-context information to guide behavior without any parameter updates. Importantly, sensory observations in held-out tasks are entirely novel to the model. We focus on an offline setting, where the in-context dataset $\mathcal{D}$ is generated by a fixed exploratory policy. The model learns to use this experience to construct a policy in this previously unseen task—effectively performing reinforcement learning in-context.
Structure of task suites
Our first task distribution is a $5\times 5$ gridworld in which the reward location is fixed but hidden from the agent during exploration. This setting is loosely inspired by the Morris water maze, a behavioral task used to study how animals leverage memory to navigate unknown environments [Vorhees and Williams, 2006]. Across tasks in this distribution, the transition function $\mathcal{T}$ is fixed, while the state encoding $\mathcal{S}$ and reward location $\mathcal{R}$ vary (Fig. 1B). At test time, the model is deployed in a gridworld with novel sensory observations and a new reward location.
@@
Our second task distribution consists of tree-structured mazes, which introduce hierarchical state transitions and sparse rewards [Rosenberg et al., 2021]. The meta-training set consists of binary trees generated with a branching probability of 0.9, producing diverse but related maze structures across tasks (Fig. 1C). The action space consists of four options: stay, move to the parent node, or move to either child node. Invalid transitions (e.g., from the root to a parent) result in the agent remaining in place. As in the gridworld setting, the state encodings $\mathcal{S}$ and reward locations $\mathcal{R}$ vary across tasks. At test time, the model is evaluated on a full 7-layer binary tree, consistent with Rosenberg et al. [2021], again using novel state encodings not seen during training.
In both task suites, each state is represented by a 10-dimensional Gaussian vector, with spatial correlations introduced across states. The in-context dataset $\mathcal{D}$ is generated using a random walk policy augmented with heuristics designed to mimic animal-like exploration patterns (Appendix A). Further task design details are provided in Appendix A. Together, these two tasks allow us to analyze model behavior in both a spatially regular setting (gridworld) and a structured, branching environment (tree maze). We note that tree mazes are setting where rodents display rapid learning [Rosenberg et al., 2021]. More broadly, branching environment structures pose a challenge for standard sequence models—such as language models trained on text—due to its non-Euclidean, hierarchical layout and lack of surface-level structure [Bogin and Berant, 2022, Ruiz and Nachum, 2021].
Model architecture and selection
@@
Our base architecture is a causal, GPT2-style transformer with 3 layers and 512-dimensional embeddings. To study how experiential memory supports decision-making, we provide the context memory $\mathcal{D}$ before the query token $s_{query}$. This contrasts with Lee et al. [2023], where the query appears first. This ordering supports an interpretation in which previous experiences are stored as cached key–value memories that are retrieved at query time. A useful analogy is the autoregressive interpretation of linear attention transformers [Katharopoulos et al., 2020]. To efficiently train the model with this token ordering, we implement custom attention masks during training (Appendix B). We meta-train five models from different random seeds and select the best-performing one for analysis in the main results. Additional details on model architecture, training procedures, and parameter sweeps are provided in Appendix B.
Several alternative modeling choices are possible, including using uncorrelated stimuli or increasing the number of transformer layers. We evaluate these variations in Appendix C, comparing their outcomes to the main results. None of these modifications meaningfully affected the results; further discussion is provided in Appendix C.
@@
Figure 2: Transformers can rapidly learn and plan in new tasks. A. Average max-normalized return in two held-out gridworld environments as a function of context length. For each length, $20$ query states are sampled at random and the test horizon is $15$. B. As in (A), but return is plotted against the number of rewards experienced in-context, and we collect results over $50$ held-out environments. Each point represents the results from one environment; averages are shown as line plots. Colors denote the meta-learned transformer (blue), tabular Q-learning (green), and DQN (pink). C. Example of shortcut behavior in a held-out gridworld. The model is shown a circuitous trajectory (orange dashed line) from start (red dot) to goal (red star), but executes a more efficient path (blue). D. As in (A), but for held-out tree maze environments, where the test horizon is $100$. E. As in (B), but for held-out tree maze environments. F. As in (D,E), but shown only for context length $800$ and subdivided by query type: states seen before reward (Pre-$\star$), after reward (Post-$\star$), or never seen in context (Novel; not used in B,E). Model types use the same color scheme as in (E). Error bars show 95% confidence intervals over the $50$ sample environments.
@@ 4 Results
4.1 Transformers learn a RL strategy to rapidly solve planning tasks relevant to natural cognition.
We evaluate the agent’s ability to act in a goal-directed manner from a query state, using an in-context dataset generated by an exploratory policy in a new environment with novel sensory observations.
@@
Gridworld  We begin by evaluating the agent’s behavior in held-out gridworld environments. We evaluate performance from query states that were observed in-context and located at least $6$ steps from the goal. We plot return as a function of context length (Fig. 2A; Appendix D). If no eligible query states meet the selection criteria, return is recorded as zero. In many cases, the meta-learned agent navigates directly to the reward after a single exposure, mirroring one-shot learning reported in rodents navigating water mazes [Steele and Morris, 1999]. To summarize performance across test environments, we plot return as a function of the number of rewards experienced in-context (Fig. 2B, blue). As expected from the single-trial plots in Fig. 2A, the agent achieves near-maximal performance after just one reward exposure, with only minor improvements thereafter.
@@
We next compare the meta-learned model to standard reinforcement learning methods. Specifically, we train a tabular Q-learning agent and a deep Q-network (DQN) on each test environment. The tabular agent serves as a useful control, isolating performance when the challenge of learning representations is abstracted away. Each Q-learning agent is trained using a replay buffer containing the same in-context dataset $\mathcal{D}$ provided to the meta-learned model. We train each agent for multiple epochs until convergence (see Appendix E for details). As before, we summarize performance by plotting return as a function of the number of rewards experienced in-context (Fig. 2B). Both Q-learning agents underperform relative to the meta-learned model, with the tabular agent slightly outperforming the DQN. The tabular agent’s advantage over the DQN highlights that representation learning adds additional difficulty in novel environments. The performance gap between the Q-learning agents and the meta-learned model is substantial but expected, reflecting the utility of learned priors in this task distribution. These findings suggest that describing the learning efficiency seen in humans and animals may require moving beyond standard reinforcement learning frameworks.
@@
Finally, we observe that the meta-learned agent discovers shortcut paths to reward. Specifically, even when the agent only observes a circuitous path in-context, it infers a policy that selects the shortest route to reward—often through previously unseen states (Fig. 2C). Quantitatively, the model selects shortcut paths in over 60% of test simulations (Appendix D), compared to 2% under a chance policy. This suggests that the agent has internalized a representation of the Euclidean geometry of the environment, a feature that we analyze more deeply in later sections.
@@
Tree Mazes We next evaluate the meta-learned agent in a tree maze environment inspired by Rosenberg et al. [2021]. As in gridworld, the agent rapidly learns the task after only a few reward exposures (Fig. 2D–E; Appendix D). We repeat the comparison with Q-learning agents in the tree maze environment (Fig. 2E). As before, learned priors enable the meta-learned agent to capture rapid learning more effectively than the Q-learning baselines. The tabular agent again outperforms the DQN, confirming that representation learning remains a core challenge in novel environments.
@@
To gain insight into the priors acquired through meta-learning, we evaluate all models at a long context length ($800$ timesteps). We stratify performance by the type of query state (i.e., start state; Fig. 2F). When the query state was experienced prior to any reward, the tabular agent performed comparably to the meta-learned model. However, when the query state appeared only after the final reward, the tabular agent underperforms. This suggests that the meta-learned model acquires a useful prior: the ability to infer inverse actions. In standard Q-learning, states encountered only along paths away from reward do not receive value propagation. Finally, we evaluate performance when the query state was never encountered during context. In this setting, the meta-learned agent performs better than both Q-learning baselines, likely due to a learned prior over action selection.
@@
Figure 3: Model representations are shaped by in-context structure learning. A. Example from a test gridworld. Query token representations are visualized for each state after projection onto the first two principal components, with outputs shown from layer 1 (left), layer 2 (middle), and layer 3 (right). Context length is 1. Points are colored by graph location; gray lines indicate true connectivity. Reward is marked with a red star. B. As in (A), but with context length 250. C. Kernel alignment between model representations and latent graph structure as a function of context length, across $100$ environments. Shaded regions indicate 95% confidence intervals; colors denote model layer. Dashed line shows baseline from raw inputs. D. As in (C), but shown only for context length 250 and including reward-ablated conditions (shaded bars). E, F. As in (A, B), but for test tree mazes. Points are colored by maze depth. G, H. As in (C, D), but for test tree mazes.
4.2 The RL strategy learned by the model relies on in-context structure learning to shape representations.
@@
Gridworld We begin by visualizing low-dimensional embeddings of model activity across gridworld states. Because the query token represents the agent’s current state, we extract its activity as the basis for representation analysis. We project the $512$-dimensional query representations into a 2D PCA space at each model layer, for both short and long context lengths (Fig. 3A–B; Appendix F). Each point is colored by the spatial location of the corresponding query state in the gridworld. With limited context, the model’s representations are disorganized and show no spatial structure (Fig. 3A). As in-context experience increases, the representations become more structured and reflect the latent geometry of the gridworld (Fig. 3B). This structure resembles predictive representation learning, but crucially, no such objective was imposed during training.
@@
To quantify this effect, we compute the kernel alignment between model representations and the latent environment structure across held-out tasks (Fig. 3C; Appendix G). Kernel alignment increases with context length, with layer 2 consistently exhibiting the strongest correspondence to latent structure. Surprisingly, representation structure is largely unaffected by the presence or absence of reward (Fig. 3D; Appendix FG). Overall, we find that in-context experience in gridworld induces geometry-aligned state representations.
@@
Tree Mazes We next ask whether in-context representation learning also emerges in agents trained on the tree maze task. We repeat the PCA projection analysis on query token representations in tree mazes (Fig. 3E–F; Appendix F). As context increases, representations organize into a bifurcating structure that roughly mirrors the maze’s hierarchical layout (Fig. 3F). Kernel alignment also increases with context length in tree mazes (Fig. 3G), but remains lower than in the gridworld task. This aligns with the visualizations in Fig. 3F, which suggest that the model learns coarse, high-level structure rather than fine-grained spatial layout. Representations in the tree maze are more strongly modulated by reward (Fig. 3H; Appendix FG).
@@
Taken together, these results indicate that the model meta-learns in-context representation strategies that vary in granularity across task domains. This provides normative support for the hypothesis that structure learning facilitates efficient decision-making, and in fact in-context structure learning has also been found in the representations of large language models [Park et al., 2024]. This observation parallels predictive map formation in the hippocampus [Stachenfeld et al., 2017]—long hypothesized as a computational scaffold for memory—and suggests that similar principles can emerge in artificial agents through meta-learning.
@@
Figure 4: As context grows, representations across environments with similar structure are increasingly aligned. A. Schematic of cross-environment alignment, based on Whittington et al. [2020]. Although sensory inputs differ, environments share latent structure, and representations of matching latent states should be similar. B. Average pairwise Pearson correlation of node representations across $100$ test gridworlds, as a function of context length. Solid lines: same-node comparisons. Dashed lines: different-node comparisons. Shading indicates 95% confidence interval. Line color denotes model layer. C. PCA visualization of representations pooled from 50 gridworlds. Points from 15 randomly selected environments are shown for clarity. D. As in (B), but for $50$ test tree mazes. E. As in (C), but for test tree mazes. F. Summary of (D) at context length 1600, averaged across layers. Cross-environment correlation scores are binned by the maze-layer of the comparison node.
4.3 Representation learning also aligns representations across environments with shared structure.
@@
In neuroscience, the hippocampal–entorhinal circuit is thought to support structure learning across contexts [Buckmaster et al., 2004, Kumaran et al., 2009, Whittington et al., 2020]. A leading hypothesis holds that the hippocampus encodes context-specific experiences, while the entorhinal cortex abstracts shared structure across environments [Whittington et al., 2020]. Do similar cross-context alignment strategies emerge in meta-learned agents?

### v3

@@ 1 Introduction
Animals can learn efficiently and rapidly adapt to new environments with minimal experience. For example, humans can infer underlying structure or learn new concepts from just a handful of examples and mice in maze tasks can identify optimal paths after only a few successful trials (meister2022learning). Standard reinforcement learning (RL) algorithms, which typically rely on incremental value updates to shape decision-making, does not capture this rapid learning behavior well (eckstein2024hybrid). One explanation is that biological agents possess useful priors shaped by evolution and experience, allowing them to generalize quickly in naturalistic settings. They also rely on episodic memory, the ability to recall specific past experiences to guide decisions in novel situations.
Here, we ask how episodic memory operates not just as storage, but as a computational substrate for rapid learning and decision-making. We train a transformer model to perform in-context reinforcement learning (lee2023supervised) on navigation tasks inspired by rodent behavior. In each new environment, the model receives exploratory trajectories as context and infers a goal-directed policy. Transformers are especially relevant not only because of their established capabilities for rapid and flexible in-context learning (dong2022survey; brown2020language; lampinen2024broader), but also because their key–value memory architecture has been linked to models of episodic memory in the brain (krotov2020large; tyulmankov2021biological; fang_2025). Understanding the reinforcement learning strategy that emerges in these models can provide new hypotheses for how memory-based computations might support flexible decision-making in new environments.
We focus on two task suites: spatially regular gridworlds and hierarchically structured tree mazes (Fig 1BC). While both require memory to support goal-directed behavior, they differ sharply in geometry: gridworlds are Euclidean and spatially continuous, whereas tree mazes are non-Euclidean and branch-structured. This contrast allows us to evaluate how learned in-context strategies generalize across structural regimes known to challenge standard sequence models– for instance, language models are known to struggle with symbolic reasoning and hierarchical generalization in tree-structured domains (bogin2022probing; ruiz2021graph; keysers2020measuring).
@@
  • We demonstrate that the model learns computations found in natural cognition. Representation learning strategies are consistent with suggested roles for hippocampus and entorhinal cortex, and memory recall patterns at decision time are consistent with hippocampal replay.
@@
  • We give descriptions for the in-context RL algorithms that emerge. We show with mechanistic analysis that the model does not use standard model-free or model-based RL methods. Instead, strategies tend to rely on intermediate computations cached in memory tokens, demonstrating how episodic memory can be used as a computational workspace.
@@
Figure 1: A transformer is trained to in-context reinforcement learn in diverse planning tasks. A. Diagram of meta-learning setup. For each task, the model is trained via supervision to predict the optimal action from a query state $s_{\text{query}}$, given memories of RL transition tuples sampled in-context. B. Illustration of three training tasks (orange) and one test task (gray) from the gridworld distribution. In each task, the underlying graph structure is fixed, but the reward location (red star) can vary. Each state is encoded as a random Gaussian vector (bottom). Importantly, test task state encodings are novel. C. As in (B), but for the tree maze distribution. The training set graph structures are drawn from probabilistically branching trees, while the test set structure is a full binary tree.
Related works
@@
In meta-RL settings, the outer learning loop shapes the weights of the network to learn an algorithm that can be deployed in-context. The in-context learning within each task occurs via activation dynamics– through memory and internal state– rather than parameter updates (beck2023survey; sandbrink2024modelling; lampinen2024broader). Early examples of this approach include RL2, which meta-trains a recurrent neural network using RL in the outer loop, such that an in-context RL strategy emerges in the inner loop (wang2016learning; duan2016rl). Subsequent work extended this approach to include explicit episodic memory mechanisms, combining RNNs with key–value memory architectures (ritter2018been; ritter2020rapid; team2023human). More recently, lee2023supervised proposed decision-pretrained transformers (DPTs), which use supervised training in the outer loop to induce in-context RL behavior in the inner loop. We adopt DPTs in this work both for practical reasons–their scalability and ease of training–and for scientific ones: the transformer architecture allows us to probe how memory-based computation supports in-context learning. This latter motivation is inspired by recent findings suggest that key–value architectures, like the transformer, offer a useful computational analogy for episodic memory (krotov2020large; ramsauer2020hopfield; tyulmankov2021biological; whittington2021relating; fang_2025; chandra2025episodic). See App. A for a more thorough discussion of related work.
@@ 2 Experimental Methods
Meta-learning procedure We adopt the decision-pretraining framework from lee2023supervised (Fig. 1A), training models via supervision to learn optimal policies from in-context experience. Each task is a Markov Decision Process defined by state encoding function $\mathcal{S}$, action space $\mathcal{A}$, transition function $\mathcal{T}$, reward function $\mathcal{R}$. For each task, the model receives an in- context dataset $\mathcal{D}$ of RL transition tuples $(s,a,s^{\prime},r)$ gathered from an exploratory policy, plus a query state $s_{\text{query}}\in\mathcal{S}$. The model is meta-trained to predict the action from an oracle policy. At test time, the model generalizes to held-out tasks with novel sensory observations using only in-context information, demonstrating in-context reinforcement learning. We focus on offline settings where $\mathcal{D}$ comes from random exploration, and more details can be found in App. B.
@@
Structure of task suites Our first task distribution is a $5\times 5$ gridworld in which the reward location is fixed but hidden from the agent. This setting is loosely inspired by the Morris water maze, a behavioral task used to study how animals use memory to navigate unknown environments (vorhees2006morris). Across tasks, $\mathcal{T}$ is fixed, while $\mathcal{S}$ and $\mathcal{R}$ vary (Fig. 1B). At test time, the model is deployed in a gridworld with novel sensory observations and a new reward location. Our second task distribution consists of tree-structured mazes, which introduce hierarchical state transitions and sparse rewards. These are settings where rodents have been shown to display rapid learning (rosenberg2021mice). The meta-training set consists of binary trees generated with some branching probability so that $\mathcal{T}$ varies across tasks (Fig. 1C). $\mathcal{S}$ and $\mathcal{R}$ also vary across tasks. The action space consists of four options: stay, move to the parent node, or move to either child node. At test time, the model is evaluated on a full 7-layer binary tree, consistent with rosenberg2021mice, again using novel state encodings not seen during training.
In both tasks, states are represented by 10-dimensional random vectors. Full task details are in App. B. Together, these two tasks allow us to analyze model behavior in spatially regular environments and branching tree structures (which may be a relevant analogy for language generation tasks).
@@
Model architecture and selection Our base architecture is a causal, GPT2-style transformer with 3 layers and 512-dimensional embeddings. We provide the context memory $\mathcal{D}$ before the query token $s_{query}$. This ordering supports an interpretation in which previous experiences are stored as cached key–value memories that are retrieved at query, or decision-making, time. Additional details on architecture and training are in App. C. We also test alternative modeling choices in App. D.
@@
Figure 2: Transformers can rapidly learn and plan in new tasks. A. Average max-normalized return in two held-out gridworld environments as a function of context length. For each context length, 2020 query states are sampled with test horizon $15$. B. As in (A), but return is plotted against the number of rewards experienced in-context and averaged over $50$ held-out environments. Blue: meta-learned transformer; Green: tabular Q-learning; Pink: DQN. C. Example of shortcut behavior in a held-out gridworld. The model experiences a circuitous trajectory (orange), but can infer a more efficient path (blue). D., E. As in (A, B), but for tree mazes and test horizon $100$. F. As in (D,E), but shown only for context length $800$ and subdivided by query type: states seen before reward (Pre-$\star$), after reward (Post-$\star$), or never seen in context (Novel; not used in E). Error bars show 95% C.I.
@@ 3 Results
3.1 Transformers learn a RL strategy to rapidly solve planning tasks.
We first evaluate the agent’s performance in new environments with novel sensory observations.
@@
Gridworld  In held-out gridworld environments, we test the agent with query states that were observed in-context and located at least $6$ steps from the goal. We plot return as a function of context length (Fig. 2A; App. E). The meta-learned agent often navigates directly to the reward after a single exposure, mirroring one-shot learning reported in rodents navigating water mazes (steele1999delay). To summarize test environment performance, we plot return as a function of the number of rewards experienced in-context (Fig. 2B, blue). As expected from Fig. 2A, the agent achieves near-maximal performance after just one exposure, with only minor improvements thereafter.
@@
We next compare the meta-learned model to standard reinforcement learning methods. Specifically, we train a tabular Q-learning agent and a deep Q-network (DQN) on each test environment. Each Q-learning agent is trained using a replay buffer containing the same in-context dataset $\mathcal{D}$ provided to the meta-learned model (App. F). We again summarize performance in Fig. 2B. The performance gap between both Q-learning agents and the meta-learned model is substantial but expected, reflecting the utility of meta-learned priors. The advantage of the tabular agent over DQN demonstrates that representation learning adds additional difficulty in novel environments. Overall, describing the learning efficiency seen in animals may require moving beyond single-task RL frameworks.
@@
Finally, we observe that the meta-learned agent discovers shortcut paths to reward. Even when the agent only observes a circuitous path in-context, it infers a policy that selects the shortest route to reward—often through previously unseen states (Fig. 2C). Quantitatively, the model selects shortcut paths in over 60% of test simulations (App. E). This suggests that the agent has internalized the Euclidean geometry of the environment, a feature that we analyze more deeply in later sections.
@@
Tree Mazes We next evaluate the meta-learned agent in test tree mazes, where the agent rapidly learns the task after only a few reward exposures (Fig. 2DE; App. E). As before, the meta-learned agent captures rapid learning more effectively than Q-learning baselines (Fig. 2E). The tabular agent again outperforms the DQN, confirming that representation learning remains a core challenge.
@@
To gain insight into the priors acquired through meta-learning, we evaluate all models at a long context length ($800$ timesteps). We stratify performance by the type of query state (Fig. 2F). When the query state had already been seen prior to any reward, the tabular agent performed comparably to the meta-learned model. However, when the query state was seen only after the final reward, the tabular agent underperforms. In standard Q-learning, states encountered only along paths away from reward do not receive value propagation. This suggests that the meta-learned model acquires a useful prior: the ability to infer inverse actions. Finally, we evaluate performance when the query state was never encountered during context. In this setting, the meta-learned agent performs better than both Q-learning baselines, likely due to a learned prior over action selection.
@@
Figure 3: Model representations are shaped by in-context structure learning. A. An example test gridworld. Query token representations are visualized for each state after projection onto the first two principal components, for layer 1 (left), layer 2 (middle), and layer 3 (right). Context length is 1. Points are colored by graph location; gray lines indicate true connectivity. Reward is marked with a red star. B. As in (A), but with context length 250. C. Kernel alignment between model representations and latent graph structure as a function of context length, across $100$ environments. Shading shows 95% C.I.; colors denote model layer. Dashed line shows baseline from raw inputs. D. As in (C), but for context length 250, with reward ablation (shaded bars). E, F. As in (A, B), but for test tree mazes. Points are colored by maze depth. G, H. As in (C, D), but for test tree mazes.
3.2 RL strategy shapes representations via in-context structure learning.
@@
Gridworld We begin by visualizing low-dimensional embeddings of model activity across gridworld states. Because the query token represents the agent’s current state, we extract its activity as the basis for representation analysis. We project the $512$-dimensional query representations into a 2D PCA space at each model layer, for both short and long context lengths (Fig. 3A–B; App. G). With limited context, the model’s representations are disorganized and show no spatial structure (Fig. 3A). As in-context experience increases, the representations become more structured and reflect the latent geometry of the gridworld (Fig. 3B). This structure resembles predictive representation learning, but crucially, no such objective was imposed during training.
@@
To quantify this, we compute the kernel alignment between model representations and the latent environment structure across held-out tasks (Fig. 3C; App. H). Kernel alignment increases with context length, with layer 2 consistently exhibiting the strongest correspondence to latent structure. Surprisingly, representation structure is largely unaffected by the presence of reward (Fig. 3D; App. GH). Overall, we find that in-context experience induces geometry-aligned state representations.
@@
Tree Mazes Does in-context representation learning also emerge in agents trained on the tree maze task? We repeat the PCA projection analysis on query token representations in tree mazes (Fig. 3E–F; App. G). As context increases, representations organize into a bifurcating structure that roughly mirrors the maze’s hierarchical layout (Fig. 3F), which suggests that the model learns coarse, high-level structure rather than fine-grained spatial layout. Consistent with this intuition, kernel alignment also increases with context length in tree mazes (Fig. 3G), but remains lower than in the gridworld task. Representations in the tree maze are more strongly modulated by reward (Fig. 3H; App. GH).
@@
Taken together, these results indicate that the model meta-learns in-context representation strategies that vary in granularity across task domains. This provides normative support for the hypothesis that structure learning facilitates efficient decision-making, and in fact in-context structure learning has also been found in the representations of large language models (park2024iclr). This observation parallels predictive map formation in the hippocampus (stachenfeld2017hippocampus)—long hypothesized as a computational scaffold for memory—and suggests that similar principles can emerge in artificial agents through meta-learning.
@@
Figure 4: As context grows, representations across environments with similar structure are aligned. A. Diagram of cross-environment alignment (whittington2020tolman). Although sensory inputs differ, environments share latent structure, and representations of matching latent states should be similar. B. Average pairwise Pearson correlation coefficient of node representations across $100$ gridworld environments, as a function of context length. Solid lines: same-node comparisons. Dashed lines: different-node comparisons. Shading shows 95% C.I.. Line color denotes model layer. C. PCA visualization of representations pooled from 15 randomly selected visualizations. D. As in (B), but for $50$ tree mazes. E. As in (C), but for tree mazes. F. Summary of (D) at context length 1600, averaged across layers. X-axis denotes the maze-layer of the comparison node.
3.3 Representations are reused across environments with shared structure.
@@
In neuroscience, the hippocampal–entorhinal circuit is thought to support structure learning across contexts (buckmaster2004entorhinal; kumaran2009tracking; whittington2020tolman). A leading hypothesis holds that the hippocampus encodes context-specific experiences, while the entorhinal cortex abstracts shared structure across environments (whittington2020tolman). Do similar cross-context alignment strategies emerge in meta-learned agents?